OpenStudy (jessie13):

1: Solve -5 + x/3 = -1. 12 -12 2 -2 2: Solve the following equation, and check your solution. Be sure to show all work in order to receive full credit. Solve 24 - 3x = -27. 3: Solve the following equation, and check your solution. Be sure to show all work in order to receive full credit. Solve -17 + n/5 = 33.

OpenStudy (jessie13):

@sophiaholder

OpenStudy (sophiaholder):

Okay one second

OpenStudy (jessie13):

k

OpenStudy (sophiaholder):

1. C Okay I am working on 2 and 3. I keep getting them wrong.

OpenStudy (jessie13):

oh okay thanks ill wait!

OpenStudy (nincompoop):

uh

OpenStudy (sophiaholder):

2. 24-24 - 3x = -27-24 -3x/-3 = -51/-3 x = 17

OpenStudy (sophiaholder):

I can't work them out while typing them I have to do it on paper. Just a tiny mistake.

OpenStudy (jessie13):

k

OpenStudy (nincompoop):

\(24 +27 -3x +3x= -27 +27 +3x \)

OpenStudy (jessie13):

for which one? @nincompoop

OpenStudy (nincompoop):

2

OpenStudy (jessie13):

for which one? @nincompoop

OpenStudy (sophiaholder):

Okay 3 is 250

OpenStudy (jessie13):

how would i show all my work? @sophiaholder @nincompoop

OpenStudy (sophiaholder):

-17+17 + n/5 = 33+17 n/5*5 = 50*5 m = 250

OpenStudy (sophiaholder):

There you go! *= MUltiply

OpenStudy (sophiaholder):

Sorry that took so long. I am usually pretty fast at algebra. It is one of my fave subjects.

OpenStudy (jessie13):

its okay!

OpenStudy (jessie13):

how would i show all my work? @nincompoop

OpenStudy (nincompoop):

type it here

OpenStudy (sophiaholder):

1. is 2. Do you need to show your work for that one?

OpenStudy (jessie13):

no just the two last ones!@sophiaholder

OpenStudy (jessie13):

What do u mean?? @nincompoop

OpenStudy (sophiaholder):

1. 2 2. 24-24 - 3x = -27-24 -3x/-3 = -51/-3 x = 17 3. -17+17 + n/5 = 33+17 n/5*5 = 50*5 m = 250

OpenStudy (nincompoop):

nice dogs

OpenStudy (sophiaholder):

lol

OpenStudy (jessie13):

okay thanks!! @sophiaholder

OpenStudy (sophiaholder):

No problem! :3

OpenStudy (nincompoop):

why m? the variable for 3 is n not m

OpenStudy (sophiaholder):

It doesn't really matter. That just happened to be the letter I used. She can change it

OpenStudy (nincompoop):

OH OKAY